OpenStudy (anonymous):

Find the equation of a line that is perpendicular to y=3x-8 and intersects this line at its y intercept.

OpenStudy (mathmale):

If the slope of one line is m, and you want the slope of a 2nd line which is perpendicular to the first, find it by taking the negative reciprocal of the slope, m, of the 1st line. You are given y=3x-8. What's the slope of this line? What's the negative of this slope? What's the negative reciprocal of this slope? Your answer is the slope of the new (perpendicular) line.

OpenStudy (phi):

mathmale was helping. the first step is to find the slope of your equation y=3x-8 Do you know the slope ?

OpenStudy (anonymous):

no i don't know the slope, I think thats whats confusing me the most.

OpenStudy (phi):

oh, I'm thinking you were supposed to have learned it by now. but if you see y = m x + b where m and b are numbers (notice this matches your equation) the number in front of the x (it's "name" is m) is the slope and the "b" is the y-intercept in your equation, : y=3x-8 what is the slope ? i.e. what is the number in front of the x ?

OpenStudy (anonymous):

oh ok so the slope is 3 and the y intercept is -8.

OpenStudy (phi):

yes. that is good. Here is a graph of this line (notice it goes through the y-axis at x=0, y=-8) and it goes over 1 step and up 3 steps (that is what a slope of 3 means)

OpenStudy (phi):

They want a line that is perpendicular. As math male was saying the slope of a line that is perpendicular (makes a right angle) is found by 1. "flipping" the slope 2. multiplying by -1 examples: slope is ½ , perpendicular slope is 2/1 * -1 or 2*-1 or -2 or -2 : think of it as -2/1 , now flip it,1/-2 or - ½ , now multiply by -1 to get ½ what is 3 when you flip it, and multiply by -1 ?

OpenStudy (anonymous):

-1/3

OpenStudy (phi):

so the perpendicular line, starting with the "generic" equation y= mx + b we know m is -⅓ \[ y= -\frac{1}{3} x+b\] we know the point (0,-8) is supposed to be on this line (because they tell us intersects this line at its y intercept i.e. at (0,-8) so put in 0 for x and -8 for y

OpenStudy (phi):

you should get this: \[ -8= -\frac{1}{3}\cdot 0 + b\] now simplify the right side

OpenStudy (phi):

what is -⅓ * 0 ?

OpenStudy (anonymous):

0

OpenStudy (phi):

so now you have -8= 0+b what is 0+b simplify to?

OpenStudy (phi):

adding 0 to something does not change it. 0+1= 1 0+2= 2 0+b = b

OpenStudy (anonymous):

so now its -8=b

OpenStudy (phi):

which means b is -8 so we put in -8 for b in the equation \[ y= -\frac{1}{3} x+b\]

OpenStudy (anonymous):

Ohh ok thank u so much this makes so much more sense now.

OpenStudy (phi):

you should get \[ y= -\frac{1}{3}x-8\] Here is what the lines look like.

OpenStudy (anonymous):

Now i get it thank u for helping me.